OpenStudy (anonymous):

A rectangular table is five times as long as it is wide. If an area is 80ft^2, find the length and width of the table. The width of the table is _ft The length of the table is _ft Please help.

OpenStudy (lalaly):

area of rectangle = length * width 80 = 5w*w 80=5w^2 w^2=16 w=4 so width is 4 ft length is 5*4=20 ft

OpenStudy (anonymous):

Thanks for showing the work. Big help

OpenStudy (lalaly):

:)